### v1

# MMAU: A Holistic Benchmark of Agent Capabilities Across Diverse Domains

Abstract
Recent advances in large language models (LLMs) have increased the demand for comprehensive benchmarks to evaluate their capabilities as human-like agents. Existing benchmarks, while useful, often focus on specific application scenarios, emphasizing task completion but failing to dissect the underlying skills that drive these outcomes. This lack of granularity makes it difficult to deeply discern where failures stem from. Additionally, setting up these environments requires considerable effort, and issues of unreliability and reproducibility sometimes arise, especially in interactive tasks. To address these limitations, we introduce the Massive Multitask Agent Understanding (MMAU) benchmark, featuring comprehensive offline tasks that eliminate the need for complex environment setups. It evaluates models across five domains, including Tool-use, Directed Acyclic Graph (DAG) QA, Data Science and Machine Learning coding, Contest-level programming and Mathematics, and covers five essential capabilities: Understanding, Reasoning, Planning, Problem-solving, and Self-correction. With a total of 20 meticulously designed tasks encompassing over 3K distinct prompts, MMAU provides a comprehensive framework for evaluating the strengths and limitations of LLM agents. By testing 18 representative models on MMAU, we provide deep and insightful analyses. Ultimately, MMAU not only sheds light on the capabilities and limitations of LLM agents but also enhances the interpretability of their performance. Datasets and evaluation scripts of MMAU are released at https://github.com/apple/axlearn/docs/research/mmau.
Figure 1: Evaluation results across different models on MMAU. For clarity, this figure includes only a selection of representative models. The domain-centric, capability-centric, and overall evaluation results are aggregated from all 20 tasks in MMAU. For detailed per-task evaluations, please refer to Appendix C.
1 Introduction
Figure 2: Overview of MMAU. MMAU is designed to provide both capability-centric evaluation (top) and domain-centric evaluation (bottom). It includes over 3K distinct prompts spanning 64 subjects and 5 domains. To evaluate the fundamental capabilities of LLM agents in a disentangled manner, we carefully designed 20 tasks aimed at decomposing these capabilities and assessing performance. Note: For clear visualization, the data examples and prompts here are simplified to illustrate an intuitive example. For the exact data examples and prompts, please refer to the Appendix D B.
Recent advancements in the field of AI have been marked by significant progresses in the development of LLMs. Particularly, one promising direction along this evolution is the ability of LLMs to perform as human-like agents [1], i.e., understand complex contexts, reason and plan with complicated logic [2, 3, 4, 5], make decisions, and utilize tools effectively [6, 7, 8]. Consequently, the need for comprehensive benchmarks that evaluate LLMs as intelligent agents has become more and more important.
While existing benchmarks [9, 10, 11, 12] evaluate LLM agents by focusing on specific application scenarios and task completion, they struggle to reveal the underlying capabilities driving these outcomes. For example, as shown in Fig. 3, when an LLM encounters a complex math problem, multiple capabilities are required to solve it. By emphasizing task completion, existing benchmarks often obscure whether a failure stems from a lack of comprehension, reasoning, or calculation error. Consequently, these evaluation methods blur the distinctions between different types of failures, hindering our understanding of where the error originates from and limiting our ability to gain deeper insights into the model’s capabilities and make targeted improvements. Additionally, some tasks in existing benchmarks require considerable effort to set up the environments, making a thorough evaluation both expensive and challenging. Furthermore, we observe that tasks, especially interactive ones, are sometimes less stable and reproducible due to the stochasticity of the environment feedback during the evaluation process. This randomness can make it difficult to obtain consistent evaluation results and draw solid conclusions.
To address such limitations, we introduce the Massive Multitask Agent Understanding (MMAU) benchmark. We develop MMAU by identifing five essential capabilities: Understanding, Reasoning, Planning, Problem-solving and Self-correction across five domains: Tool-use, Directed Acyclic Graph (DAG) QA, Data Science &\&& Machine Learning coding, Contest-level programming, and Mathematics. As a result, MMAU comprises a total of 3,220 distinct prompts gathered from diverse data sources. These include our in-house human annotations for tool-use, as well as rewritten and curated prompts from open-source datasets such as CodeContest [13], Kaggle [14], and DeepMind-Math [15]. Based on this dataset, we designed 20 tasks across 64 subjects, offering a comprehensive benchmark. To avoid the complexities of environment setup and issues of unreliability, all tasks in MMAU are performed on our 3K static dataset to eliminate potential issues related to environment instability.
We comprehensively evaluate 18 models on MMAU, which include both API-based commercial models and open-source models. In addition to conventional overall comparisons and evaluations tailored to specific application scenarios, we also study the varying capabilities across different models (Figure 1). From our study utilizing MMAU, thorough analysis and insightful analysis arise (Sec. 5). We envision MMAU as a valuable benchmark that not only yields significant observations but also equips the community with deeper insights. Our contributions are summarized as:
  • We offers evaluations from both application scenarios and fundamental capabilities, providing a comprehensive framework for understanding the strengths and limitations of LLM agents.
  • The evaluation process on MMAU is straightforward and unified on a static dataset, avoiding the instability issues that may arise from interactive evaluations and thus ensuring reliable results.
  • We release our evaluation dataset and scripts, aiming to set a new standard for performance assessment in the AI landscape.
We also acknowledge that interactive evaluation is necessary. MMAU does not aim to replace them but rather to complement them by addressing the issues mentioned above. Developing and providing a more stable and easy-to-use benchmark for interactive evaluations is valuable and warrants further studies.
Figure 3: Different error types on a math problem.
2 Related Work
LLM-based Genelist Agents Researchers have proposed various generalist agent frameworks to create AI assistants that can understand and execute any instruction from users. One pioneering work is Auto-GPT [16], which uses a language model as an AI agent that can break down goals into actionable steps with the aid of auxiliary tools. Integrating language models into multi-agent collaboration systems [17] is a cutting-edge research area. Frameworks like AutoGen [18], LangChain [19], Camel [20], AGENTS [21], AutoAgents [22], and XAgent [23] have explored different approaches to enable communicative agents to collaborate autonomously, facilitate practical applications, ensure control and customization, dynamically generate specialized agents, and manage complex tasks effectively.
Agent Benchmarks The need for rigorous benchmarks also arises in response to the surge in LLM-based agents. Varying benchmarks are created to gauge agent capability on tasks inspired by real-world use cases. The Berkeley Function Calling Leaderboard [24], NexusRaven V2 Function Calling Benchmark [25], ToolBench [26], StableToolBench [27], and API-BLEND [28] seek to evaluate the capability of LLM agent to plan and perform function calls. Webshop [29], WebArena [30], Mind2Web [31], MiniWoB++ [32], and VisualWebArena [33] focus on the agent’s ability to browse and interact with a web environment. A line of benchmarks consider the universal presence of user interface (UI) and envision UI automation agents, including PixelHelp [34], MetaGUI [35], MoTIF [36], AITW [37], and OmniACT [38]. SWE-bench [39] tests agent capability to solve real-world software engineering problems.
While each benchmark tends to focus on a specific application, a generalist agent should be able to perform well on a wide range of tasks. AgentBench [12] consolidates tasks covering coding, game, and math into a single systematic benchmark. AgentBoard [11] evaluates agent capabilities under web browsing, tool use, embodied AI, and game domains. However, both benchmarks require containerized environments to run and need involved effort to implement new tasks. As a result, simply stacking tasks would lead to diminishing returns in comprehensiveness. MMAU takes a step back and considers a range of core agent capabilities. Based on the core capabilities, MMAU provides a range of tasks that are designed to produce decoupled metrics over the core capabilities.
Table 1 compares the supporting capabilities of various benchmarks. PlanBench [40], specifically designed for benchmarking planning capabilities in LLM agents, also supports disentangled evaluation of reasoning and planning. AgendaBoard [11] provides a manually labeled subset for evaluating disentangled capabilities, focusing on interactive agents, such as spatial navigation and world modeling. In contrast, MMAU evaluates more fundamental and essential capabilities for LLM agents. Overall, MMAU offers a more comprehensive evaluation of fundamental capabilities in a disentangled manner.
3 The MMAU Benchmark
Table 1: Comparison of benchmarks in evaluating core capabilities of LLM agents. “En.” and “Dis.” represent entangled and disentangled, specifically. Understand.: understanding, Reason.: reasoning, Plan.: planning, Prob.-solv.: problem-solving, Self-corr.: self-correction, MM: multimodal grounding.
Benchmarks Understand. Reason. Plan. Prob.-solv. Self-corr. MM
En. Dis. En. Dis. En. Dis. En. Dis.
AgentBench [12]
AgentBoard [11]
PlanBench [40]
MMLU [9]
MMMU [10]
MMAU
To introduce MMAU, we will start with an overview of all included capabilities 3.1. We will then provide detailed explanations of how each task was designed and how the dataset across different domains was constructed 3.2.
3.1 Capabilities in MMAU
Below, we introduce the capability definitions and the key tasks used to evaluate them. A complete task-capability mapping can be found in the Appendix 6.
Understanding is a fundamental capability required of an intelligent agent. In MMAU, we evaluate an agent’s understanding in different aspects, including: complex instruction following, user intent understanding, statistics parsing, and visual grounding.
Reasoning and Planning reflect an agent’s thought process and ability to infer logically from complex factors. Although reasoning and planning has been recognized in many works, it is often described as a general ability, compounded with other skills, which limits deep investigation. In MMAU, we address this challenge with the task planner-shift, designed to decompose reasoning and planning capabilities from other factors. Unlike standard end-to-end evaluations, planner-shift divides the solution generation into two stages. In stage 1, a planner model generates a high-level plan, providing a strategy to solve the given problem without hinting at the final solution. In stage 2, a solver model is given the original problem along with the plan to solve it. This approach isolates the planning and reasoning processes from problem-solving. To test planning and reasoning capabilities, we vary only the planner model while using the same solver model, ensuring that performance differences reflect the planning and reasoning capabilities. The task design diagram is shown in Figure 4.
Problem-solving focuses on measuring an agent’s ability to successfully implement or execute a task, assuming it has already understood and planned the strategy well. To address this, we design a task called solver-shift, similar to planner-shift, which also performs a two-stage generation. However, solver-shift keeps the planner model constant and varies only the solver models to reflect differences in problem-solving skills, as shown in Figure 4. In MMAU we use the tasks of planner-shift and solver-shift in domains of Contest-level coding 3.2.3 and Math 3.2.4.
Self-correction is another core capability for an intelligent agent. It reflects the agent’s ability to identify errors, learn from its environment and past behaviors, and correct itself to eventually overcome obstacles and achieve its task. In MMAU, we evaluate this capability by specifically designing self-correction tasks across different domains.
Figure 4: Construction of planner-shift task and solver-shift task.
3.2 Dataset Construction
The construction of MMAU encompasses both breadth and depth of data, as illustrated in Table 4. Our dataset is constructed from heterogeneous sources: 1) our in-house tool-use data, used for tasks under tool-use and DAG-QA; 2) Kaggle [14] datasets, which we rewrote to design tasks for DS &\&& ML coding; 3) CodeContest [13], used for tasks under contest-level coding; and 4) DeepMind-math [15], used for math tasks. MMAU involves curating and rewriting these data sources. In the following section, we will explain how we leveraged these source data to construct MMAU.
3.2.1 Tool-Use
We curated an in-house dataset for tool-use with conversation trajectories following the standard tool-use (a.k.a. function-calling) protocol. We select from a subset of RapidAPI Hub  [^1] https://rapidapi.com/hub functions and ask human annotators to create realistic scenarios with user queries, ground truth function calls and actual function returns from the RapidAPI endpoints. In total, our in-house tool-use dataset consists of 409 single-step (task:single-tool-use) and 258 multi-step tool-use conversations (task: multi-turn multi-tool-use). Out of the 409 single-step tool use conversations, 225 require making parallel tool calls (task: parallel-tool-use). Figure LABEL:fig:tool-use-examples shows some of these examples. We adapt this dataset for the following tasks.
Task: Tool-use Benchmarking agent tool-use following the standard protocol requires an interactive environment. To simplify the evaluation process, we instead evaluate the model’s response at each assistant turn (i.e., where a function call is expected), conditioning on the ground-truth versions of all previous user or assistant turns. For evaluation, we check if the model’s tool call matches that of the ground truth, i.e. calling the same function and the same parameters.
Task: DAG QA In this task, a user presents a set of requirements to which the LLM must respond by selecting and ordering a sequence of tool invocations from multiple choices provided. This design examines whether the model can identify the relevant tools and deduce the correct dependencies between them. The prompt enumerates the possible tool use orderings from which the LLM agent is asked to pick one, and the label is derived from the ground truth function call sequence. For example, we transform the multi-step example in Figure LABEL:fig:tool-use-examples into a DAG QA task as shown in Figure 14
Task: Tool-use Self-correction From the tool-use dataset described above, we derive two classes of errors to test the model’s self-correction ability:
Temporary error simulates a tool that is temporarily unavailable. From the ground truth messages [user queries, tool-calls and tool responses], we substitute a random temporary error (e.g. “429: too many requests”, “504: gateway timeout”) in place of the tool response.
Incorrect call simulates a previous tool call or response containing an error. We mutate the ground truth tool-call to be incorrect by changing the arguments or the function name, issue the modified call, and save the updated tool response. Given the message history of [user queries, mutated calls, updated tool responses], the model is expected to retry with the correct call.
The evaluation metric is the exact match accuracy of the function name and arguments against the ground truth.
Figure 5: A Multi-turn coding and QA example for Data science and Machine learning.
3.2.2 Data Science and Machine Learning
We leverage the Meta Kaggle Code dataset [14] and curate 28 Python notebook-style conversations, with 123 conversation turns. Each turn begins with a user request for code generation. Among all requests, 83 requests expect text-based outputs from code and 40 requests expect image outputs. Due to the open-ended nature of code generation, we created multiple-choice questions that require information from successful code execution to fully address, resulting in 207 text-based questions and 121 image-based questions. Figure 5 shows an example turn with multiple choice questions. We report QA accuracy as the main metric and vary the combination of code model and QA model to produce different evaluation settings.
Task: E2E Code Generation and QA In this setting, we aim to gauge the overall capability. the evaluated model is responsible for both code generation and QA.
Task: Code Generation and GPT-4 QA In this setting, we isolate the code generation capability of the model. After generating code from the evaluated model, we adopt a strong multimodal model (GPT-4 [41]) to serve as control and perform QA based on code execution outputs.
Task: QA from Oracle Code In this setting, we specifically focus on the textual and visual understanding proficiency of the model decoupled from code generation. We obtain oracle output by executing ground truth code implementation and then pass to the evaluated model to perform QA.
Task: DS &\&& ML Self-Correction This setting is similar to the E2E setting, however, whenever code execution fails, we use the execution error message to prompt an additional code generation turn.
3.2.3 Contest-Level Coding
For contest-level coding problems, we select 261 problems from the Valid and Test splits of the CodeContests dataset [13] which includes competitive programming problems. We adapt these 261 problems for the following tasks.
Task: E2E Standard In this task, models are challenged with a variety of coding problems. The effectiveness of the solutions is measured by executing the code against all predefined test cases [13]. All CodeContests results reported in this paper are based on pass@K (K=5) accuracy.
Task: Planner-shift and Solver-shift As introduced in Section 3.1, we use thses two tasks to extensively measure the agent’s capability in planning and problem-solving, respectively. We evaluate both of these tasks by generating K Python code solutions and verifying their pass rate.
Task: Problem Parsing Unlike the other tasks, this task does not require the model to write or execute any code. Instead, given a problem statement and associated test cases, the model is only tasked with predicting the outputs for these test cases. An example is shown in Figure 15. Our rationale is that if a model truly grasps the problem, including its complex instructions and user intent, it should be able to accurately predict the outputs based on its understanding alone. We use match accuracy as the evaluate metric.
Task: CodeContest Self-correction For the E2E standard task above, we collect the error messages of each candidate solution if it does not pass some test cases, including 4 types of errors: empty solution, compilation error, runtime error, and wrong outputs. Then, we follow the setup as what we used in Tool-use self-correction (Sec. 3.2.1) to append the error content as a feedback user message to the message list, and ask the model to try again. We still measure the pass@K metric by generating K candidates independently.
3.2.4 Mathematics
The source data in the domain of math, derived from DeepMind-Math [15], consists of 1,000 carefully curated math problems spanning 56 subjects, including calculus, geometry, statistics, and etc.
Task: E2E Standard We adhere to standard protocol by incorporating a Chain-of-Thought (CoT) [3] into the prompt to generate the answers end-to-end, and using accuracy as the evaluation metric.
Task: Planner-shift and Solver-shift As introduced in Sec. 3.1, we use the two tasks to assess the model’s planning and problem-solving abilities in a two-stage manner, avoiding confounding influences from other capabilities. The prompts used for each task can be found in the Appendix B.
Task: Comprehend+. To better isolate and assess the understanding capability without excessive interference from other skills, we have devised a new task named Comprehend+. Our hypothesis is that problems that are straightforward mathematically but complex in their descriptions rely more heavily on understanding capabilities. To test this, we first selected a subset containing only the mathematically simpler problems, and then use an LLM to create new math problems that feature more complex descriptions or harder problem statements but retain the same underlying mathematical constructs from each data sample. A rewritten math problem example is shown in Fig. 13. After curation and verification, we finalize 676 newly created problems for Comprehend+. Please refer to Appendix A.5 for details of dataset creation.
4 Evaluation
We comprehensively evaluate 18 models on MMAU. All evaluation model details are listed in Table 5. For easier reading, the main paper presents only the aggregated evaluation results. For the evaluation results over all 20 tasks, please refer to Appendix C.
4.1 Domain-centric Evaluation and Analysis
[…]
Model Tool-use DAG DS&\&&ML CodeContests Math
GPT-4o [42] 69.33 77.38 66.52 31.80 53.40
GPT-4-Turbo [43] 75.13 79.38 63.90 25.67 38.57
GPT-3.5-Turbo 66.63 28.60 35.79 10.34 25.00
Gemini-1.5-pro [44] 52.43 71.18 54.02 10.34 39.70
Gemini-1.0-pro [45] 26.32 47.45 27.87 7.66 39.40
Claude3 Opus [46] 62.39 73.61 59.45 15.33 37.40
Claude3 Sonnet [46] 51.92 57.65 52.44 10.34 26.80
Claude3 Haiku [46] 44.14 36.36 33.84 8.81 36.60
Mixtral-8x22B-v0.1 [47] N/A 72.51 31.10 9.20 50.00
Mixtral-8x7B-v0.1 [47] N/A 30.82 13.48 1.92 21.70
Mistral-7B-v0.2 [48] N/A 34.15 2.80 0.38 9.55
Phi-3-mini4K-instruct [49] N/A 23.95 0.84 2.30 21.00
Openfunctions-v2 [50] 26.53* 25.28 3.84 8.05 15.75
Hermes-2-Pro-Mistral-7B [51] 39.48 27.94 3.29 0.77 12.78
Command R [52] 28.29* 22.28 0.00 4.21 8.21
LLama2-70B [53] N/A 19.73 0.00 0.00 8.43
Llama2-13B [53] N/A 17.96 0.00 0.00 4.10
Llama2-7B [53] N/A 20.84 0.00 0.00 3.92
As shown in Table 2, there is a clear performance gap between API-based commercial models and open-source models across all evaluation domains. Among the commercial models, the GPT-4 family (including GPT-4o and GPT-4) consistently outperforms other models. In math and contest-level coding, GPT-4o demonstrates a significant advantage. Additionally, Claude3-Opus and Gemini-1.5-pro perform reasonably well. While for open-source models, a significant number do not support tool-use. Among those that do, Hermes-2-Pro-Mistral-7B demonstrates strong tool-use performance. For models that do not support tool use, Mixtral-8x22B-Instruct-v0.1 performs surprisingly well in math and DAG-QA, demonstrating its strong reasoning and planning capability. Additionally, Phi-3 performs well in math considering its model size. The Llama2 family, however, struggles with challenging coding tasks.
4.2 Capability-centric Evaluation and Analysis
As introduced in Sec.3.1, we designed tasks to decompose core capabilities from standard evaluations, allowing MMAU to offer a unique dimension of evaluation. Each capability includes tasks spanning different domains. To provide overall capability-centric evaluation results for each model, we aggregate tasks under each capability using a weighted average. Detailed task-capability mappings and the calculation method can be found in Appendix A. The overall capability-centric evaluation results are shown in Table 3.
Notably, for the capability of Understanding, GPT-4o significantly outperforms other models, demonstrating its superior capability in handling long contexts, complex user instructions, and capturing (sometimes implicit) user intents. Additionally, GPT-4, Gemini-1.5-pro, and Claude3-Opus also exhibit reasonably strong understanding capabilities. For the capabilities of Reasoning and Planning, the GPT-4 family shows the strongest performance. When examining the capability of Problem-solving, the performance gap is not significantly large. This trend suggests that when provided with "oracle" plans, solving a task may be less challenging. While models’ problem-solving capabilities vary, most can perform these tasks reasonably well, indicating that this capability may be more universally achievable among different models. On the contrary, for Self-correction, we observe a significant gap among models. Among open-source models, aside from Mixtral-8x22B, others do not seem to possess the skill to reflect on and correct their own errors effectively. These evaluation results highlight that self-correction is a critical capability needing further research and development to advance the field.
[…]
Problem-Solving Self-correct
Model w/o Tool-Use w/ Tool-Use Understanding Reasoning Planning w/o Tool-Use w/ Tool-Use
GPT-4o [42] 56.12 61.11 60.63 50.47 47.90 43.65 51.56
GPT-4-Turbo [43] 48.07 58.29 49.78 50.88 49.59 40.86 51.86
GPT-3.5-Turbo 42.84 51.83 30.78 29.38 28.27 21.23 32.38
Gemini-1.5-pro [44] 49.04 50.32 47.63 36.28 33.77 34.69 36.66
Gemini-1.0-pro [45] 52.84 42.82 37.05 39.88 37.31 15.96 10.91
Claude3 Opus [46] 49.98 54.67 49.03 44.10 38.84 38.47 44.16
Claude3 Sonnet [46] 40.47 44.79 40.22 36.57 36.92 31.01 37.05
Claude3 Haiku [46] 39.77 41.42 36.85 42.09 42.09 20.44 30.01
Mixtral-8x22B-v0.1 [47] 49.04 N/A 44.39 44.92 38.02 18.00 N/A
Mixtral-8x7B-v0.1 [47] 33.50 N/A 27.98 27.57 30.67 8.26 N/A
Mistral-7B-v0.2 [48] 21.87 N/A 9.01 21.45 21.22 1.93 N/A
Phi-3-mini4K-instruct [49] 22.94 N/A 14.92 28.33 27.45 2.04 N/A
Openfunctions-v2 [50] 20.89 29.43* 11.20 23.30 24.76 2.61 22.77*
Hermes-2-Pro-Mistral-7B [51] 29.26 33.12 16.54 25.95 24.96 2.04 2.99
Command R [52] 22.17 31.30* 19.47 22.72 22.95 0.35 18.84*
LLama2-70B [53] 24.23 N/A 6.32 15.30 14.55 0.34 N/A
Llama2-13B [53] 12.86 N/A 2.69 13.39 13.15 0 N/A
Llama2-7B [53] 15.92 N/A 2.38 14.25 13.26 0 N/A
5 Analysis and discussion
[…]
Do different capabilities present varying levels of difficulty for models to achieve? Our evaluation results reveal that different capabilities indeed present varying levels of difficulty for models to achieve. As what we mentioned in Sec. 4.2, problem-solving capabilities exhibit a smaller performance gap among models, suggesting that problem-solving is a more universally achievable capability across different models. However, self-correction capabilities present a significant challenge, with a notable performance gap observed among models, and many open-source models lacking effective self-correction skills. These findings suggest that while some capabilities like problem-solving are more readily attained by current models, others, such as self-correction and planning pose greater challenges and are vital areas for future advancements in the field.
Do balanced capabilities indicate the path to a generalist agent? From Figure 1, we observe that strong models, such as the GPT-4 family, exhibit balanced performance across all capabilities, demonstrating their robust and versatile nature. This balance indicates that improvements in one area likely enhance performance in others, highlighting a high correlation and interdependence among these capabilities. Conversely, models that perform poorly in one capability tend to struggle across the board, suggesting underlying weaknesses in their architecture or training strategy.
Are larger models always better? Another interesting finding arises when comparing the MistralAI families and Llama-2 families. For the MistralAI models, we consistently observe performance gains with increasing model size across domains. However, this trend does not apply to the Llama-2 families. In code-related domains (DS &\&& ML, CodeContest), all size variants of Llama-2 perform poorly. Surprisingly, in the DAG-QA domain, the Llama-2-7B model performs better than its larger counterparts. This observation is consistent with findings from AgentBench [12], further validating that training strategies and model architectures also influence the scaling law.
6 Conclusion
In this paper, we introduce the Massive Multitask Agent Understanding (MMAU) benchmark. By evaluating models based on both application scenarios and fundamental capabilities, MMAU provides a comprehensive and in-depth test bed for reliable and thorough studies. By designing 20 tasks to decompose capabilities beyond standard evaluation benchmarks, MMAU offers more granular insights into the strengths and limitations of these models.
Limitations and future work. The current scope of MMAU, while broad, does however not encompass all possible domains relevant to LLM agents, such as interactive environments which are also critical yet challenging. Future iterations of MMAU should aim to include interactive tasks to provide a more holistic evaluation. This expansion will require the development of reliable, stable, and user-friendly interactive environments. Moreover, as we expand to include more domains, it will be essential to incorporate additional capabilities such as retrieving, memorizing, sequential decision-making, etc. Our current approach to capability decomposition, though insightful, still faces challenges in disentangling compound capabilities. Future research should focus on developing more effective methods for decomposing and evaluating these capabilities to further refine the benchmark.
Ethics and Societal Impacts. Research on LLM agents must consider potential ethical concerns and negative societal impacts. The MMAU benchmark aims to provide a thorough and transparent evaluation framework, but it is crucial to ensure that these evaluations do not inadvertently reinforce biases or propagate harmful content. We are also careful in detecting and mitigating any personally identifiable information or offensive content within our datasets and prompts.
Acknowledgement
We acknowledge and thank Mark Lee who helped with our code review and release.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix D
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Code, data, and instructions needed to reproduce the main experimental results are included in the supplemental material.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Appendix D
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix A
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix A
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Experiment and dataset details
The key statistics of the MMAU dataset are presented in Tab. 4. A list of all evaluated models (commercial and open source) is provided in Tab. 5.
Table 4: The key statistics of MMAU.
Source data subjects task prompts turns Answer type
Tool-use In-house Sports, Health, Location, etc single-tool. 409 N/A function-call
multi-tool. 258 N/A function-call
DAG QA 695 N/A multi-choice
self-correct 282 N/A function-call
CodeContest CodeContest [13] alg., datastr., etc standard 261 N/A execution
problem parsing multi-choice
PlannerShift execution
Solvershift execution
ML&\&&DS Kaggle[14] DS, ML, Visual textual QA 207 N/A multi-choice
code generation 123 N/A execution
visual QA 121 N/A multi-choice
Math DM-math[15] calculus, geometry… standard 1K N/A math solution
PlannerShift
SolverShift
synthetic Comprehend+ 676
Table 5: Evaluation models in MMAU.
API-based Commertial Models Open-source Models
GPT-4o-2024-05-13 Mixtral-8x22B-Instruct-v0.1
GPT-4-turbo-2024-04-09 Mixtral-8x7B-Instruct-v0.1
GPT-3.5-turbo-0125 Mistral-7B-Instruct-v0.2
Gemini-1.5-pro-preview-0409 Phi-3-mini-4k-instruct
Gemini-1.0-pro gorilla-openfunctions-v2
Claude-3-opus-20240229 Hermes-2-Pro-Mistral-7B
Claude-3-sonnet-20240229 c4ai-command-r-v01
Claude-3-haiku-20240307 c4ai-command-r-rplus
Llama-2-70b-chat-hf
Llama-2-13b-chat-hf
Llama-2-7b-chat-hf
For our evaluations the open source models have been hosted via VLLM  [^6] https://github.com/vllm-project/vllm on up to 8 NVidia A100 or H100 GPUs per model depending on the size. We have used the checkpoints from Huggingface for the open source models.
In order to have deterministic prediction we use a temperature of 0 and greedy search for all models and tasks.
A.1 Tool-use
Data construction protocol: 1) The user sends a query to the agent model along with a list of potential functions including a description of their purposes and parameters. 2) The agent responds with either natural language or appropriate function use. 3) In case of function-call, the functions are invoked according to the agent’s instructions, either by the user or directly by the agent, and the result is submitted back to the agent model. 4) The agent can then conclude with the given information or continue the conversation with follow-up questions or additional function calls.
Tool-use evaluation details: To compare the predicted and the ground truth parameter values, we perform string normalization including stripping punctuation, white spaces and converting to lower case. In some cases where the parameter value can have open-ended, semantically equivalent forms, we define example-specific match rules based on regular expressions to accommodate valid alternatives.
A.2 DAG QA
Data construction protocol: The query includes a description of the task (to choose the appropriate plan), a list of potential functions (including a description of their purposes and parameters), an enumeration of all possible plans (all possible sequences in which to execute the tools), and the task input (what the task is that the plan is expected to solve). The task is formulated as a multiple choice task, where at the end of the query, the agent is asked to end the reply with which plan it chooses. At evaluation, the chosen plan is then extracted from the output by searching for a string match of the requested response format. An example prompt is illustrated in Figure 14.
The reasoning and planning benchmarks differ only in the prompt, where for reasoning, the agent is requested to "elaborate on the thought process and reasoning", while for planning, the agent is requested to "be concise with a response in the format Chosen Plan: N".
A.3 Self-correction tool-use
To encourage models to retry, we prepend the following system message to the user’s first turn message:
Figure 6: retry-message
Kaggle: To represent all results on the same scale and reduce confusion, when QA model is not multimodal, we report its performance as $\mathrm{ratio\_of\_text\_based\_questions}*\mathrm{accuracy\_on\_text\_based\_questions}=(207/328)*\mathrm{accuracy\_on\_text\_based\_questions}$. We set the temperature to 0 for both code generation and QA.
A.4 Contest-level Coding
We select 261 valid problems from the Valid and Test splits of the CodeContests dataset [13], each contains a number of test cases, including public tests, private tests, and generated tests.
E2E Standard. The agent is asked to solve the problem by generating Python code solution. The prompt contains a detailed description of the CodeContests problem with public test examples, as well as basic instructions on code formatting following the baseline prompt design in [54]. The generated code solution will be compiled and executed on all test cases, and is considered correct only if all test cases are passed. Following related works, we report the pass@5 score, which is the percentage of problems solved by generating 5 solutions per problem. We set temperature as 0.3 for all models.
Problem Parsing. We design a ProblemParsing task for each CodeContests problem, which provides the agent with the problem description and public test examples, and asks the agent to directly infer the desired output for an unseen test case. To make the unseen test case relatively easy for parsing and computation, we select the shorted test case from the union of private test cases and generated test cases. We adopt a chain-of-thought prompt to help the LLM agent understand the problem (see Appendix B for detailed prompts). We then compare the answer with the groundtruth output, and derive the accuracy over the entire dataset. For reproducibility, we set temperature as 0 for all models.
Planner-shift. As introduced in 3.2.3, we divide the code generation process for CodeContests problems into two steps: generating a plan (i.e., planner) and generating code solutions (i.e., solver). We evaluate the planning capability of LLM agents by fixing the solver to be the same strong model. For each LLM model, we generate 1 plan with temperature 0, and use gpt-4-0125-preview as the solver to generate 5 code solutions (with temperature 0.3) and calculate the percentage of solved problems.
Solver-shift. SolverShift complements PlannerShift by freezing the planner and evaluating the problem-solving capability of various LLM models. Given 1 plan generated by gpt-4-0125-preview, we let LLM agents generate 5 code solutions and calculate the percentage of solved problems.
Self-correction / Retry. We also evaluate the self-correct capability of LLM agents on the CodeContests problems. For the Regular task above, we collect the error messages of each candidate solution if it does not pass some test cases, including 4 types of errors: empty solution, compilation error, runtime error, and wrong outputs. For the problems where all the 5 candidates fail to pass, we will append the error content as a feedback user message in the message list, and ask the LLM to try again. To encourage the agent to retry, we also prepend an additional system instruction to the first user message before describing the problem. The evaluation metric is also pass@5.
A.5 Math
Given a set of seed math problems $S$, we first employ a LLM as the judger $M_{J}$ to determine the difficulty level for each problem, resulting in a labeled math problem set $S^{\prime}:\{s_{i}^{d}\}$. We then select a subset containing only the mathematically simpler problems, i.e., $S^{\prime\prime}:\{s_{i}^{d},d\leq\theta\}$. As a result we have a curated “simple” math problem set $S^{\prime\prime}$ (672 problems). Next, we use another LLM $M_{w}$ (GPT-4 [41]) to create new math problems that feature more complex descriptions or harder problem statements but retain the same underlying mathematical constructs for each data sample $S^{\prime\prime\prime}:\{M_{w}(s_{i}^{d}),d\leq\theta\}$. A rewrite math problem example is shown in Fig. 13. Finally, to ensure that each rewritten problem in $S^{\prime\prime\prime}$ remains valid and has the same ground truth answer as $S$, we perform a data curation stage. As a result, we finalize the problem set $\mathcal{S^{*}}$ for the Comprehend+ task.
Table 6: Mapping of capabilities and tasks.
Capability Tool-use DAG DS&ML CodeContests Math
Problem-solving Tool Execution - Code Generation and GPT-4 QA Solver-shift Solver-shift
Understanding - - QA from Oracle Code Problem Parsing Comprehend+
Reasoning - DAG QA w/ Reasoning - Planner-shift Planner-shift
Planning - DAG QA w/o Reasoning - Planner-shift Planner-shift
Self-correct Self-correction - Self-correction Self-correction -
Appendix B Prompt templates
B.1 Tool-use
For tool use the concrete prompts are generated by the model to transform the tool definitions as well as the predicted tool calls and tool results into the concrete prompt. Therefore, we don’t report concrete prompt here. Samples from the dataset are show in Figs. 10, 11, and 12.
B.2 DAG
Listing 1: DAG Reasoning Prompt
Choose the appropriate plan with its associated set of tools to accomplish a specific task.
Each tools functionality is using Open API schema notation.
"[A B]" indicates that tools A and B can operate concurrently without any interdependencies.
"A -> B" indicates that tool A must be executed prior to tool B, as Bs operation is contingent upon the output from A.
"B -> A" indicates that tool B must be executed prior to tool A, as As operation is contingent upon the output from B.
"[A B] -> C" indicates that tools A and B will be executed in parallel initially, followed by tool C, which requires the outputs from both A and B.
Below is a compilation of all available tools, described in JSON format:
{tool_descriptions}
Here are all possible plans:
{possible_plans}
Task Input:
{task_input}
Feel free to elaborate on your thought process and reasoning.
Please strictly end the reply with "Chosen Plan: N", where "N" must be an integer number of the selected plan.
Do not change the answer format to other words and MUST include Chosen Plan:
Listing 2: DAG Planning Prompt
Choose the appropriate plan with its associated set of tools to accomplish a specific task.
Each tools functionality is using Open API schema notation.
"[A B]" indicates that tools A and B can operate concurrently without any interdependencies.
"A -> B" indicates that tool A must be executed prior to tool B, as Bs operation is contingent upon the output from A.
"B -> A" indicates that tool B must be executed prior to tool A, as As operation is contingent upon the output from B.
"[A B] -> C" indicates that tools A and B will be executed in parallel initially, followed by tool C, which requires the outputs from both A and B.
Below is a compilation of all available tools, described in JSON format:
{tool_descriptions}
Here are all possible plans:
{possible_plans}
Task Input:
{task_input}
Be concise with a response in the format "Chosen Plan: N", where "N" represents the number of the selected plan.
B.3 Data Science and Machine Learning
Listing 3: DS & ML Code Generation Prompt
You are a helpful code assistant that will help users analyze data and answering multiple choice questions based on the data analysis.
You will be provided with a Jupyter notebook file (.ipynb) containing code cells and markdown cells.
A list of data files used in the notebook will also be provided.
The notebook would contain a code interpreter style conversation between a user and an assistant.
User would be asking questions, follow up with additional requests, and providing clarification for requests.
Assistant would be analyzing data with code interpreter, providing insights into data based on analysis results and diagrams, and responding to user inquiries.
Your task is to address the users request in the last markdown cell of the notebook AND answer the multiple choice questions based on the data analysis.
You will first append an appropriate code cell to the notebook to address the users request.
You will then be provided with the execution results of the code cell. Based on the results, you will answer the multiple choice questions.
For example, given
Data Files:
[’../input/pokemon/pokemon.csv’]
Notebook:
[…]
User Request:
[…]
You will respond with markdown cell and code cell to address the users request:
[…]
Then, you will be provided with the execution results of the code cell and multiple choice questions based on the results:
Execution Output: # the images would be attached in the same order as the image placeholders
[{"data": {"image/png": "<image_1>"}, "execution_count": 1, "metadata": {}, "output_type": "display_data"}]
Questions:
[…]
You will then provide the answers to the multiple choice questions:
["B", "B", "C", "B"]
You MUST exactly follow nbformat structure when generating markdown and code cells.
You are only allowed to use the following math and visualization packages: numpy, pandas, scipy, scikit-learn, sympy, statsmodels, matplotlib, seaborn, plotly, wordcloud. Python standard library is also allowed. Make sure to import the necessary packages in the code cell.
You MUST answer in a list of uppercase single letter strings for multiple choice questions.
Now, given the following notebook and user request, respond with a list containing a markdown cell and a code cell to address the users request:
Data Files:
{data_files}
Notebook:
{notebook}
User Request:
{user_request}
You MUST exactly follow nbformat structure when generating markdown and code cells.
Respond ONLY with the cells in json format. DO NOT say anything else or put it in a ‘‘‘json‘‘‘ code block.
You are only allowed to use the following math and visualization packages: numpy, pandas, scipy, scikit-learn, sympy, statsmodels, matplotlib, seaborn, plotly, wordcloud. Python standard libraries are also allowed. Make sure to import the necessary packages in the code cell.
Listing 4: DS & ML QA Prompt
Given the following execution output and multiple choice questions, respond with the answers to the questions:
Execution Output:
{execution_output}
Questions:
{questions}
You MUST answer in a list of uppercase single letter strings for multiple choice questions.
Listing 5: DS & ML Execution Feedback Prompt
Your code execution failed. Please check the error message below and try again:
{error_message}
You MUST exactly follow nbformat structure when generating markdown and code cells.
Respond ONLY with the cells in json format. DO NOT say anything else or put it in a ‘‘‘json‘‘‘ code block.
You are only allowed to use the following math and visualization packages: numpy, pandas, scipy, scikit-learn, sympy, statsmodels, matplotlib, seaborn, plotly, wordcloud. Python standard libraries are also allowed. Make sure to import the necessary packages in the code cell.
B.4 Contest-Level Coding
Listing 6: CodeContests E2E Standard Prompt
You are an AI with advanced code generation capabilities.
When you encounter a specific problem (labeled #PROBLEM#), your goal is to produce a valid python code that correctly solves the problem.
Make sure to fully address the problem goals following the rules and constraints.
The code should be robust and general. It should output correct answers under any valid inputs, not only just the example inputs given in the problem description.
#PROBLEM#:
{description}
The code output must follow this structure:
def f1(...):
return ...
def f2(...):
return ...
if __name__ == "__main__":
The code should read the input using the input()’ method. Make sure to properly parse the input, according to the problem description.
The output should be printed without additional words using the print()’ method.
#ANSWER#:
‘‘‘python
Listing 7: CodeContests Problem Parsing Prompt
You are an AI with advanced comprehension capabilities, akin to that of humans. This enables you to understand complex instructions, discern the users intent, and apply logical reasoning across a variety of scenarios effectively.
When you encounter a specific problem (labeled #PROBLEM#) along with a relevant example input (referred to as #TEST CASE#), your goal is to deduce the correct answer (#ANSWER#).
Please adhere to the following format:
First, express your #THOUGHTS# to reflect your understanding of the problem and test case. Keep your #THOUGHTS# concise, not exceeding 5 sentences.
Then, provide your #ANSWER# to the problem.
For example, your output can be:
#THROUGHTS#:
The problem is about ... The test cases are ...
#ANSWER#:
1 2
3 4
Here is the problem.
#PROBLEM#:
{description}
#TEST CASE#:
{inputs}
Now, please provide the #THOUGHTS# and #ANSWER# sections. Do not include additional explanations or reasoning processes.
#THROUGHTS#:
Listing 8: CodeContests Planner-shift Prompt
You are an AI with advanced code understanding and planning capabilities.
When you encounter a specific problem (labeled #PROBLEM#), your goal is to devise a structured and executable plan to solve the problem.
Your plan should clearly outline the logical steps needed to address the problem, breaking down complex processes into manageable actions.
Its crucial that each step is straightforward enough to be easily translated into executable code or functions by a Language Model (LLM).
Keep the plan simple and focused, avoiding unnecessary complexity to ensure ease of implementation.
#PROBLEM#:
{description}
#GENERATED PLAN#:
Listing 9: CodeContests Solver-shift Prompt
You are an AI with advanced code generation and instruction following capabilities.
When you encounter a specific problem (labeled #PROBLEM#) and a solving plan (labeled #PLAN#), your goal is to produce a valid python code that correctly solves the problem.
Make sure to fully address the problem goals following the rules and constraints. Refer to the plan to generate the code.
The code should be robust and general. It should output correct answers under any valid inputs, not only just the example inputs given in the problem description.
#PROBLEM#:
{description}
#PLAN#:
{plan}
guidelines:
- Generate only code, without any additional explanations or comments.
- Make sure to include all the necessary module imports, properly initialize the variables, and address the problem constraints.
- The code needs to be self-contained, and executable as-is.
The code output must follow this structure:
def f1(...):
return ...
def f2(...):
return ...
if __name__ == "__main__":
The code should read the input using the input()’ method. Make sure to properly parse the input, according to the problem description.
The output should be printed without additional words using the print()’ method.
answer:
‘‘‘python
B.5 Mathematics
Listing 10: Math Generation Prompt
You are an AI with advanced comprehension capabilities, akin to that of humans. This enables you to understand complex instructions, discern the users intent, and apply logical reasoning across a variety of scenarios effectively.
When you encounter a specific problem (labeled #PROBLEM#), your goal is to deduce the correct answer (#ANSWER#).
Please adhere to the following format:
First, express your #THOUGHTS# to reflect your understanding of the problem. Keep your #THOUGHTS# concise, not exceeding 5 sentences.
Then, provide your #ANSWER# to the problem.
For example, your output can be:
#THROUGHTS#:
Lets think step by step ...
#ANSWER#:
12
Here is the problem.
#PROBLEM#:
{description}
Now, please provide the #THOUGHTS# and #ANSWER# sections. Do not include additional explanations or reasoning processes under #ANSWER# section.
#THROUGHTS#: Lets think step by step:
#ANSWER#:
Listing 11: Math Evaluation Prompt
As a math expert, you will be provided with three items: a #Question#, an #Answer#, and the #Ground Truth#.
Your task is to determine whether the #Answer# matches the #Ground Truth#.
You need to considering mathmetical theorems when verifing the answer. For example, \$(c + 9)(c - 4)\$ and ’(c - 4)*(c + 9)’should be
the same expressions based on factorization theorems.
If they align, respond with correct’. If they do not, respond with wrong’.
Ensure that your #Response# is the final line and consists solely of the word correct or wrong’, without any additional commentary or explanation.
#Question#:
{question}
#Answer#:
{answer}
#Ground Truth#:
{ground_truth}
#Response#:
Listing 12: Math Planning Prompt
You are a math planner.
Given a problem description, your goal is to devise a structured and executable plan to solve the problem.
Your plan should clearly outline the logical steps needed to address the problem, breaking down complex processes into manageable actions.
Its crucial that each step is straightforward enough to be easily translated into executable calculation or functions by a Language Model (LLM).
Keep the plan simple and focused, avoiding unnecessary complexity to ensure ease of implementation.
Problem description:
{description}
Generated plan:
Appendix C More evaluation results
For domain-centric evaluation, we chose tasks that rely solely on the evaluation model to solve them end-to-end and adhere to the standard setup used in the field. This ensures that the domain-centric results accurately reflect the most standardized evaluation outcomes, making it easier to compare performance across different benchmarks. Specifically, we use the evalution results of E2E standard task as domain evaluation results for Contest-level coding, DS&\&&ML, Math and DAG-QA. For Tool-use, we report the weighted average of single-, parallel-, and multi-turn multi-tool-use tasks. The results are shown in Table 2.
Table 7: Evaluation resutls on math tasks. stand.: standard E2E, underst.: comprehend+, solv.-shift: solver-shift, plan.-shift: planner-shift.
Commertial Models stand. compre. solv.-shift plan.-shift OS Models stand. compre. solv.-shift plan.-shift
GPT-4o 53.4 51.48 61.2 45.2 Mixtral-8x22B 50 42.46 60.02 40.20
GPT-4-turbo 38.57 36.72 50.32 46.60 Mixtral-8x7B 21.70 27.12 43.46 28.00
GPT-3.5-turbo 25.00 20.41 45.80 32.80 Mistral-7B 9.55 25.00 27.12 15.63
Gemini-1.5-pro 39.70 35.21 56.60 25.81 Phi-3 21.00 21.00 34.23 31.30
Gemini-1.0-pro 39.40 33.13 66.67 41.67 openfunctions-v2 15.75 13.31 25.00 23.49
Claude-3-opus 37.40 36.09 56.00 36.80 Hermes-2-Pro 12.78 12.95 39.62 25.73
Claude-3-sonnet 26.80 25.44 42.90 32.30 Llama-2-70b 8.43 8.29 32.10 11.40
Claude-3-haiku 36.60 27.51 44.50 49.65 Llama-2-13b 4.10 3.85 19.84 3.85
Command-R 8.21 10.00 33.33 23.70 Llama-2-7b 3.92 2.23 24.69 25.73
Detailed tool use results are reported in Table 8.
Table 8: Detailed results for the tool use datasets. Owing to incorrect filtering, the distractor datasets contain two additional samples. They will be removed in follow-up versions.
Dataset Single-step Parallel
Multi-step
Double
Multi-step
Triple
Single-step
w. Distractors
Parallel
w. Distractors
Multi-step
Double
w. Distractors
Multi-step
Triple
w. Distractors
# Samples 184 225 298 258 185 225 300 258
Command R+ 0.739 0.662 - - 0.708 0.644 - -
Command R 0.717 0.627 - - 0.654 0.609 - -
Claude 3 Haiku 0.837 0.053 0.513 0.415 0.768 0.040 0.433 0.357
Claude 3 Opus 0.886 0.436 0.708 0.504 0.854 0.387 0.663 0.535
Claude 3 Sonnet 0.826 0.302 0.581 0.419 0.811 0.169 0.463 0.380
Gemini-1.0-pro 0.750 0.004 0.252 0.155 0.708 0.004 0.197 0.151
Gemini-1.5-flash 0.783 0.191 0.534 0.407 0.784 0.187 0.467 0.391
Gemini-1.5-pro 0.772 0.440 0.530 0.415 0.735 0.471 0.473 0.376
Openfunctions-v2 0.875 0.422 - - 0.832 0.373 - -
GPT-3.5-Turbo 0.918 0.667 0.634 0.523 0.881 0.609 0.620 0.457
GPT-4-Turbo 0.918 0.853 0.715 0.585 0.881 0.818 0.653 0.531
GPT-4o-2024-05-13 0.913 0.822 0.607 0.523 0.881 0.809 0.543 0.465
Hermes-2-Pro-Mistral-7B 0.750 0.556 0.245 0.174 0.746 0.498 0.157 0.128
For Command R, Command R+, and Hermes-2-Pro-Mistral-7B we used the tool use prompts as reported in the official model repositories  [^7] https://huggingface.co/CohereForAI/c4ai-command-r-v01  [^8] https://huggingface.co/CohereForAI/c4ai-command-r-plus  [^9] https://huggingface.co/NousResearch/Hermes-2-Pro-Mistral-7B. These prompts for Command R and Command R+ do not support multi-step tool calls, so we could not evaluate Command R and Command R+ on the double and triple datasets.
The Claude models as well as the Gemini and GPT models support tool use through dedicated APIs. These APIs have been used for the evaluation here.
Detailed Reulst on Data Science and Machine Learning
Table 9: Detailed evaluation results on Data Science and Machine Learning. The bolded and underlined numbers indicate the 1st and 2nd highest performances in each category.
Model E2E GPT4 QA Oracle Execution Retry
GPT-4o [42] 66.52 64.27 88.60 70.79
GPT-4-Turbo [43] 63.90 63.90 85.67 68.05
GPT-3.5-Turbo 35.79 58.78 52.13 37.44
Gemini-1.5-pro [44] 54.02 55.55 83.23 58.90
Gemini-1.0-pro [45] 27.87 45.12 52.93 28.66
Claude3 Opus [46] 59.45 61.65 83.48 65.12
Claude3 Sonnet [46] 52.44 56.71 75.37 54.33
Claude3 Haiku [46] 33.84 48.48 59.39 36.04
Mixtral-8x22B-v0.1 [47] 31.10 45.43 58.05 31.65
Mixtral-8x7B-v0.1 [47] 13.48 26.46 48.90 13.90
Mistral-7B-v0.2 [48] 2.80 19.94 12.40 3.17
Phi-3-mini4K-instruct [49] 0.84 4.34 0.00 3.35
Openfunctions-v2 [50] 3.84 17.07 12.74 4.02
Hermes-2-Pro-Mistral-7B [51] 3.29 17.93 23.02 3.35
Command R [52] 0.00 0.00 48.88 0.00
LLama2-70B [53] 0.00 19.51 0.00 0.00
Llama2-13B [53] 0.00 0.00 0.00 0.00
Llama2-7B [53] 0.00 0.00 0.00 0.00
Table 9 presents the evaluation results on Data Science and Machine Learning tasks, assessing E2E code generation and QA, code generation with GPT-4 QA, QA from oracle code execution, and DS & ML self-correction.
GPT-4o achieves the highest accuracy across all settings: E2E (66.52%), Execution Only + GPT-4 QA (64.27%), QA from Oracle Execution (88.60%), and E2E + Online Retry (70.79%). GPT-4-Turbo follows closely with 63.90%, 63.90%, 85.67%, and 68.05% respectively. These results highlight their robust code generation and QA capabilities.
Gemini-1.5-pro and Claude3 Opus show strong performance in QA from Oracle Execution with 83.23% and 83.48%, indicating good comprehension and QA skills, though their code generation is less effective compared to GPT-4 models. Mixtral-8x22B-v0.1 has the overall best performance among the evaluated open-source models.
Detailed Results on Contest-level Coding.
Table 10: Detailed evaluation results on Contest-level Coding. The bolded and underlined numbers indicate the 1st and 2nd highest performances in each category.
Model E2E Standard Problem Parsing PlannerShift SolverShift Retry
GPT-4o [42] 31.80 49.04 24.13 26.44 9.55
GPT-4-Turbo [43] 25.67 38.31 18.01 19.54 6.7
GPT-3.5-Turbo 10.34 30.65 17.60 11.49 0.85
Gemini-1.5-pro [44] 10.34 34.87 16.09 11.88 4.27
Gemini-1.0-pro [45] 7.66 27.20 19.92 9.58 0
Claude3 Opus [46] 15.33 39.08 21.07 12.26 4.98
Claude3 Sonnet [46] 10.34 34.10 16.48 10.73 1.71
Claude3 Haiku [46] 8.81 32.57 22.99 10.73 0.84
Mixtral-8x22B-v0.1 [47] 9.20 32.18 15.32 11.49 0.84
Mixtral-8x7B-v0.1 [47] 1.92 22.22 20.31 4.21 1.17
Mistral-7B-v0.2 [48] 0.38 8.05 21.84 4.21 0.38
Phi-3-mini4K-instruct [49] 2.30 18.01 24.52 3.07 0.39
Openfunctions-v2 [50] 8.05 3.83 19.16 9.96 0.83
Hermes-2-Pro-Mistral-7B [51] 0.77 17.62 23.37 3.83 0.39
Command R [52] 4.21 6.90 18.00 7.28 0.80
LLama2-70B [53] 0 9.20 22.61 0 0.77
Llama2-13B [53] 0 1.53 21.07 2.30 0
Llama2-7B [53] 0 5.75 20.31 2.30 0
Table 10 shows detailed results in the Contest-level Coding domain. As a challenging dataset, we find that the overall E2E Standard scores are relatively low across models, while GPT-4o and GPT-4-Turbo achieve much better results than remaining models by a large margin. GPT-4o has leading performance over all types of tasks, especially Problem Parsing and Retry, showing its remarkable capability of understanding and self-correct.
With a separation of capabilities, we can better analyze the differences and gaps among existing models. For example, we observe significant performance variation across different models on the Problem Parsing task, suggesting the intellectual challenge of understanding the contest problems due to the complicated nature of the problem description. On the other hand, we emphasize that the design of E2E Standard, SolverShift and Retry tasks require the model to generate code, while Problem Parsing and PlannerShift eliminate the actual code generation step and mainly focus on understanding/reasoning/planning capabilities. Such a fine-grained model evaluation can reflect more detailed differences between models. We can see that although some open-sourced models fail to solve most E2E Standard problems, they still possess good understanding and planning capabilities on the complicated contest-level problems (e.g. Phi-3-mini4K-instruct, Hermes-2-Pro-Mistral-7B).
By comparing the E2E Standard and Planner-shift or Solver-shift, we can see that most models have boosted performance with a separate strong solver or planner. However, the pass rate of GPT-4o and GPT-4-Turbo drops on Planner-shift and Solver-shift compared to their E2E Standard performance. This may be caused by the selection of GPT-4-0125 as the fixed solver/planner, which itself has 22.61% score on E2E Standard, stronger than most models while weaker than GPT-4o and GPT-4-Turbo.
Moreover, we find that for Contest-level Coding, having a plain step-by-step plan does not bring as significant improvement as in the math domain, as such plans may not be sufficient to solve the algorithmic challenges in the domain. A more powerful planning and reasoning structure may be adopted to better solve the Contest-level problems, as shown in [54]. But the focus of this paper is to identify and reveal the capabilities of existing LLMs, which advancing the capabilities remains as a research problem for future investigation.
Appendix D Data examples
Figure 7: Retry example where the assistant needs to recognize a temporary error and retry the call.
Figure 8: Self-correction example where the assistant needs to correct an erroneous call given explicit feedback from the tool.
Figure 9: Self-correction example where the assistant needs to recognize that the prior call, despite succeeding, did not match the user’s request, and re-issues the corrected call.
Figure 10: Example of single tool-use
Figure 11: Example of parallel tool-use
Figure 12: Example of multi-tool use
Figure 13: Example from MathComprehend+.
Figure 14: A DAG-QA example.
Figure 15: An example of CodeContests “ProblemParsing” task to measure the agent’s understanding capability.
Appendix E Datasheet for Datasets
We follow the Datasheet for Datasets framework to answer the important questions considering this dataset.
E.1 Motivation
For what purpose was the dataset created?    See Section 1.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?    Please refer to our author list for details.
Who funded the creation of the dataset?    The dataset incurred no monetary cost other than the time and effort of the authors.
E.2 Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?    See Section 3.2.
How many instances are there in total (of each type, if appropriate)?    See Table 4.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?    The full dataset is provided.
Is there a label or target associated with each instance?    See Section 3.2 for the definitions of each data type and the associated target.
Is any information missing from individual instances?    No.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?    No.
Are there recommended data splits?    The dataset is entirely intended for evaluation purpose.
Are there any errors, sources of noise, or redundancies in the dataset?    We have carefully reviewed each data sample to ensure the reliability of the benchmark. We are committed to actively improving and enriching the benchmark based on community feedback.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?    Yes, the dataset is self-contained.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?    No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?    No.
Does the dataset identify any subpopulations (e.g., by age, gender)?    No.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset?    No.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals race or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?    No.
E.3 Collection Process
How was the data associated with each instance acquired?    See Section 3.2.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?    See Section 3.2.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?    N/A.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?    The authors volunteered to perform the necessary manual data curation.
Over what timeframe was the data collected?    From April 2024 to Jun 2024.
Were any ethical review processes conducted (e.g., by an institutional review board)?    We carefully reviewed our data curation procedure. While we believe the source datasets permit our use herein, we do not make any representations as to the ethical review process of the original dataset. Please follow the terms of such datasets for further consideration.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?    See Section 3.2.
Were the individuals in question notified about the data collection?    N/A.
Did the individuals in question consent to the collection and use of their data?    N/A.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?    N/A.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted?    N/A.
E.4 Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?    See Section 3.2.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?    See Section 3.2.
Is the software that was used to preprocess/clean/label the data available?    N/A.
E.5 Uses
Has the dataset been used for any tasks already?    See Section 3.2.
Is there a repository that links to any or all papers or systems that use the dataset?    We will host the benchmark data within the Hugging Face community, which has convenient tracking features on the reference and usage of the dataset.
What (other) tasks could the dataset be used for?    The dataset can be generally customized and used for evaluating the capabilities of large language models and large multimodal models.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?    No.
Are there tasks for which the dataset should not be used?    N/A.
E.6 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?    The dataset will be fully open-sourced.
How will the dataset will be distributed (e.g., tarball on the website, API, GitHub)?    The dataset will be made available as a Hugging Face dataset.
When will the dataset be distributed?    Upon publishing the manuscript.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?    No.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances?    No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?    No.
E.7 Maintenance
Who will be supporting/hosting/maintaining the dataset?    The authors will be responsible for supporting/hosting/maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?    Please refer to our author list for details.
Is there an erratum?    No.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?    Yes.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., was the individuals in question were told that their data would be retained for a fixed period of time and then deleted)?    N/A.
Will older versions of the dataset continue to be supported/hosted/maintained?    Yes.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?    Yes. We welcome contributions via the Hugging Face community.